# Effortless, Simulation-Efficient Bayesian Inference using Tabular Foundation Models

### v1

@@ Appendix D Additional experiments
D.1 Unconditional density estimation on the UCI datasets
Here, we apply TabPFN on some classical unconditional density estimation benchmark tasks from the UCI repository [51] as used in several other works [38, 39, 90]. Note that while we still perform density estimation here, it does not involve posterior distributions. Thus, we do not refer to this approach as NPE-PF. Specifically, we consider the Gas, Power, Hepmass, and Miniboone datasets. These tabular datasets range in dimensionality from 6 to 43 features and contain between $3.1\cdot 10^{4}$ and over 1 million samples. We here investigate the unconditional density estimation performance of TabPFN in the low sample regime in comparison to a neural spline flow (NSF) [39]. As in previous works, we compute the negative log-likelihood (NLL) under a held-out test set. While we evaluate on the full test sets, we only use $10^{3}$, $10^{4}$, or (if applicable) $10^{5}$ samples for training. For the two larger settings, we use partitioning (Sec. B.3) with 10 clusters when estimating densities with TabPFN.
Figure D-1: Results for unconditional density estimation on the UCI datasets. Negative-log-likelihood (NLL) for TabPFN (with 1 and 10 clusters) and the neural spline flow (NSF) across the different UCI datasets. Dots indicate averages, and bars show standard deviation over five independent runs.
For $10^{2}$ and $10^{3}$ samples, TabPFN achieves a smaller negative log-likelihood on all four datasets (Fig. D-1). Similarly, for the $10^{4}$ samples, TabPFN outperforms NSF on all but the power dataset. Interestingly, for the lower dimensional datasets, clustering further reduces the NLL despite having access to the same total number of samples. On the other hand, for the higher dimensional datasets, clustering increases the NLL. For $10^{5}$ samples (or $3.1\cdot 10^{4}$ samples for the Miniboone dataset), clustering improves the NLL over any $10^{4}$ sample setting because it allows TabPFN to access a larger number of samples. However, NSF performs better than or equal to TabPFN on all datasets except the Gas dataset at $10^{5}$ samples.
@@ D.2 SBI benchmark with more baseline methods
We present results in terms of C2ST and SBC metrics (Fig. D-2, extending the results from the main text (Fig. 2). For SBC, we use the Error of Diagonal (EoD) to quantify the deviation from perfect calibration. NRE performs similarly or worse on all tasks, while NPE (Ensemble) matches standard NPE in terms of C2ST and achieves the best calibration of all baselines. NPE-PF is comparable to NPE (Ensemble) in terms of calibration on most tasks, while often outperforming it in predictive accuracy and simulation efficiency. Calibration plots (Fig. D-3) corroborate the EoD-based calibration summary with fine-grained information.
These results underscore that NPE-PF not only achieves strong predictive performance but also provides well-calibrated posteriors that match or exceed ensembles. While we do not directly compare NPE-PF with the work of Delaunoy et al. [33], who present a SBI method for small simulation budgets using Bayesian neural networks, their method reports modest gains over ensembles, whereas NPE-PF consistently outperforms them—further emphasizing the effectiveness of our approach with a limited simulation budget.
Figure D-2: Extended SBI benchmark results for amortized NPE-PF. (a) Extension of Fig. 2a with more baseline methods, namely NRE and NPE (Ensemble). (b) Mean absolute error between the calibration curves and the diagonal (EoD), with 0 indicating perfect calibration. Full calibration curves in Fig. D-3.
@@
D.3 Embedding networks for high-dimensional data
We evaluate the performance of NPE-PF-Infomax on high-dimensional simulation-based inference tasks, specifically focusing on a spatial SIR model [16] and an extended Lotka-Volterra task [92]. In the case of the spatial SIR task, where the true posterior is intractable, we rely on indirect metrics such as negative log-likelihood (NLL), TARP [93], and SBC for evaluation. NPE-PF-Infomax leverages compressed embeddings learned via an infomax objective to distill key information from high-dimensional observations, which are then used as inputs to NPE-PF.
@@ D.3 Embedding networks for high-dimensional data
Our results (Fig. D-4) indicate that such embedding networks—whether pre-trained or jointly trained—can be used to extend the applicability of NPE-PF to high-dimensional observations. However, in most cases, this approach does not outperform an end-to-end trained NPE baseline. Interestingly, TabPFN without an embedding network (when applicable) often achieves better performance, suggesting that the compression step, independent of NPE-PF, may be suboptimal. One possible explanation is that TabPFN, having been trained on causally structured data, can exploit underlying dependencies to improve inference accuracy. In contrast, compressing such highly structured data into low-dimensional summaries may discard critical information, limiting the effectiveness of inference.
@@
Figure D-4: High-dimensional data with embedding nets. (a) Visualization of observations in the spatial SIR task alongside posterior predictive samples obtained using NPE-PF-Infomax. (b) Performance evaluation across high-dimensional variants of the spatial SIR task over 5 independent runs, reported in terms of the negative log-likelihood (NLL) of the true parameter and the area off the diagonal in TARP and SBC calibration analyses. (c) An instance of the Lotka-Volterra task with an extended time series (300 dimensions), evaluated using NPE-PF and NPE-PF-Infomax.
D.4 Comparison of autoregressive and ratio-based density evaluation
@@
In Sec. 2.4, we present a method for fast density evaluation based on density ratios, requiring only a single forward pass through the TabPFN classifier. While this approach offers a substantial computational advantage, it is inherently approximate, raising the question of how it compares to the autoregressive density evaluation. To answer this question, we compare the two methods on two tasks—the 2D Two Moons tasks and the 10D Gaussian Linear task—across varying simulation budgets. In each case, we evaluate the density of posterior samples from the trained model given an observation, using 5 random seeds. To construct the dataset for the ratio estimator, we draw $5,000$ samples each from the posterior of interest and the uniform base distribution, which is the default setting.
@@
Figure D-5: Comparison of density evaluation approaches on Two Moons and Gaussian Linear task. (a) Correlation between density estimates from the autoregressive and ratio-based approaches for two observations using $10^{4}$ simulations. (b) Pearson correlation between the two density estimates across varying simulation budgets for both tasks. (c) Evaluation time (log scale) for an increasing number of density evaluations using an Nvidia H100 GPU, showing the crossover point where the ratio-based approach becomes more efficient.
@@
We observe a strong correlation between autoregressive and ratio-based density evaluations for the Two Moons task and a somewhat weaker correlation for the higher-dimensional Gaussian Linear task (Fig. D-5a for a simulation budget of $10^{4}$). This trend persists across simulation budgets (Fig. D-5b). In terms of runtime, the ratio-based method has an initial cost due to posterior sampling, making it slower for the first evaluation. However, once samples are obtained, subsequent evaluations are significantly faster—an advantage that becomes very noticeable as the number of density evaluations increases (Fig. D-5c; note the logarithmic y-axis).
@@
Therefore, in scenarios that require repeated density evaluations for a single observation—such as rejection sampling in TSNPE-PF—the ratio-based approach offers a highly favorable trade-off: It provides sufficient accuracy at a fraction of the computational cost.
D.5 Varying feature and noise distributions
@@
Here, we investigate the robustness of NPE-PF to different types of noise. To evaluate the robustness in such cases, we construct a variant of the Gaussian linear task from Lueckmann et al. [24] with non-Gaussian features and noise
@@
where $s$ is the scale parameter. We vary the feature dimension, the scale of feature and noise distributions, and the number of training simulations, and compare the performance of NPE-PF against NPE.
We first evaluate performance via C2ST across all combinations of feature and noise distributions for $s_{\text{features}}=s_{\text{noise}}=1.0$, $\text{dim}({\bm{\theta}})=5$, and $10^{3}$ simulations. The heatmap shows that NPE-PF outperforms NPE consistently, with while cells indicating training failure primarily for NPE (Fig. D-6a). For the example of Laplace feature and Cauchy noise distributions, we provide results as a function of simulation budget, confirming the robustness of NPE-PF across distributional shifts (Fig. D-6b). Notably, while NPE achieves performance drops on several non-Gaussian configurations, NPE-PF maintains stable performance.
Figure D-6: Robustness to feature and noise distributions. (a) Heatmaps showing C2ST performance for all combinations of feature and noise distributions with $\text{dim}({\bm{\theta}})=5$, $s=1.0$, and $10^{3}$ training simulations. White cells indicate failed training. (b) C2ST across training set sizes for an example combination of distributions. These panels are zoom-ins from (a), highlighting the stability of NPE-PF under varying distributional assumptions.
These results suggest that, although TabPFN is pre-trained using priors based on structural causal models (SCMs) with uniform or Gaussian root noise, nonlinear transformations in the SCM induce rich marginal distributions and allow generalization to a wide range of distributions. NPE-PF inherits this robustness, and its performance is largely invariant to different feature or noise distributions. Failures of NPE, especially with Cauchy features, are caused by instability in z-scoring, which NPE-PF avoids through default preprocessing such as the Yeo–Johnson transform [94]. Thus, standard NPE performance could probably be improved by applying appropriate transformations. Because NPE-PF takes advantage of the automatic preprocessing performed within TabPFN, it provides reliable performance without manual preprocessing, making it a robust and user-friendly solution in practical settings.
D.6 Order of autoregressive sampling
To sample from multi-dimensional (conditional) distributions with TabPFN, we use it in an autoregressive manner. An important question is whether the order in which we sample the dimensions matters for, e.g., the quality of the inferred posterior distributions. To investigate this question, we rerun NPE-PF on the benchmark tasks from Sec. 3.1, but permute the order in which we sample the parameter dimensions. Two of the seven benchmark tasks are permutation invariant by design (Gaussian Mixture and Gaussian Linear), and we replace these with tasks where a distribution is constructed autoregressively. The first is a simple nonlinear task given by

### v2

@@ Appendix D Additional experiments
D.1 Additional tasks
Here we extend the SBI benchmark experiments (Sec. 3.1) with additional inference tasks from other scientific fields or interesting synthetic tasks used in previous work.
Specifically, the additional tasks include simulators from physics (Weinberg, Stellar Streams), computer science (M/G/1 queue) [16], as well as “structured” synthetic tasks (Tree, HMM) [21]. Out of these, the stellar streams simulator stands out as particularly computationally demanding (>30>30 min. per simulation per CPU core; with two-dimensional parameters, and 199 dimensional data). For this reason, we investigate the performance only up to $10^{4}$ simulations.
The tasks in Hermans et al. [16] do not come with a ground truth posterior. We therefore evaluate them only via the average log posterior density (Fig.  D-1a). Overall, the results on these additional tasks again demonstrate that, for a small simulation budget, NPE-PFN significantly outperforms NPE. Specifically, NPE-PFN strongly outperforms NPE on tasks with sparse conditional dependencies, i.e., Tree, HMM, Streams, in which it is one to three orders of magnitudes more simulation-efficient than NPE. The stellar streams task is a good example use-case, where this efficiency is particularly beneficial: Attaining 100 simulations is feasible on a consumer-grade CPU (10 cores) in 5 hours, and NPE-PFN achieves a comparable performance to NPE with $10^{4}$ simulations, which would require 20 days. In addition, we investigated the calibration of NPE-PFN for which we observed a similar trend (Fig.  D-1b on a subset of tasks). Finally, we performed a direct comparison to reference posteriors where available. We also extended the HMM task to 50 parameter and data dimensions as an additional test for high-dimensional yet “structured” parameter spaces (Fig. D-1c). Here, NPE-PFN generally outperformed NPE.
The HMM example specifically illustrates how NPE (or, more generally, conditional density estimation) fundamentally struggles with dimensionality, independent of the complexity of the simulation process. One approach to address this challenge is to incorporate suitable inductive biases or constraints (such as known conditional dependencies) into the estimation network architecture [21, 30, 92]. The inductive bias from TabPFN as used in NPE-PFN seems to be especially beneficial in such cases. TabPFN was trained exclusively on synthetic datasets generated from random structural causal models (SCMs), which by construction exhibit (sparse) conditional (in-)dependencies. We therefore hypothesize that pretraining on SCMs enables TabPFN to automatically detect and exploit such (in-)dependencies directly from data. A detailed investigation of these mechanisms would be an interesting direction for future work.
Figure D-1: Results on additional tasks. (a) Performance under the average negative log likelihood metric (NLL) across five new tasks for NPE and NPE-PFN. (b) Calibration error on a subset of the new tasks. (c) Performance in C2ST against reference posteriors on tasks for which reference are available. We also investigate a different parameterization of the HMM tasks by increasing dimensionality. Note that the simulation budget of the Streams tasks only goes up to $10^{4}$ simulations.
@@ D.2 SBI benchmark with more baseline methods
We present results in terms of C2ST and SBC metrics (Fig. D-2, extending the results from the main text (Fig. 2). For SBC, we use the Error of Diagonal (EoD) to quantify the deviation from perfect calibration. NRE performs similarly or worse on all tasks, while NPE (Ensemble) matches standard NPE in terms of C2ST and achieves the best calibration of all baselines. NPE-PFN is comparable to NPE (Ensemble) in terms of calibration on most tasks, while often outperforming it in predictive accuracy and simulation efficiency. Calibration plots (Fig. D-3) corroborate the EoD-based calibration summary with fine-grained information.
These results underscore that NPE-PFN not only achieves strong predictive performance but also provides well-calibrated posteriors that match or exceed ensembles. While we do not directly compare NPE-PFN with the work of Delaunoy et al. [36], who present a SBI method for small simulation budgets using Bayesian neural networks, their method reports modest gains over ensembles, whereas NPE-PFN consistently outperforms them—further emphasizing the effectiveness of our approach with a limited simulation budget.
Throughout this manuscript, the comparisons of NPE-PFN with baseline methods are performed using the sbi library’s default hyperparameters [55]. Here, we additionally perform hyperparameter optimization of our main baseline method NPE. For each task and simulation budget, we conduct a random search for flow (flow type, number of flow layers, dimensionality of hidden flow layers) and optimization (batch size, learning rate) hyperparameters. In each setting, we limit the computing time to ten hours. For seven tasks and four simulation budgets, this results in a total computing time of 280 hours.
Compared to NPE with the default settings, NPE (Sweep) achieves better performance in some tasks, particularly with smaller simulation budgets (Fig. D-4). This is due to the shorter training times with $10^{2}$ and $10^{3}$ simulations, enabling the random search to cover large parts of the search space. For $10^{4}$ and $10^{5}$ simulations, no or only minor improvements can be observed. Further improvements for larger budgets would require more computing time and better hyperparameter optimization algorithms, such as Bayesian optimization. Despite the substantial computing time per setting, the optimized NPE (Sweep) still lags behind NPE-PFN in terms of performance for most tasks with smaller budgets (Fig. D-4). Overall, these results again demonstrate NPE-PFN’s strong default performance, providing an easy-to-use method that is competitive with, or even superior to, baseline methods where intensive hyperparameter optimization was performed. This strong default performance is particularly beneficial in the sequential setting, where hyperparameter optimization poses significant challenges due to long inference times, changing training conditions across rounds, and the difficulty of selecting suitable metrics for optimization.
Figure D-2: Extended SBI benchmark results for amortized NPE-PFN. (a) Extension of Fig. 2a with more baseline methods, namely NRE and NPE (Ensemble). (b) Mean absolute error between the calibration curves and the diagonal (EoD), with 0 indicating perfect calibration. Full calibration curves in Fig. D-3.
@@
Figure D-4: SBI benchmark results with NPE hyperparameter optimization. C2ST for NPE-PFN, NPE with default settings, and NPE (Sweep), where hyperparameter optimization was performed. Specifically flow and optimization hyperparameters are optimized separately for each benchmark task and simulation budget using ten hours of random search.
D.3 Inference speed
While NPE-PFN is completely training-free, the autoregressive use of TabPFN requires computational effort during inference (Fig. 2b). The inference speed of NPE-PFN depends on three parameters: 1) the number of simulations passed as the in-context dataset; 2) the dimensionality of the observation space (i.e., the number of features); and 3) the dimensionality of the parameter space (i.e., the number of parameters).
Here, we provide a thorough examination of the inference speed of NPE-PFN across different simulation budgets and varying dimensionalities of the parameter and observation spaces (Tab. D-2). All results were obtained using an Nvidia A100 GPU. In general, as the value of any of the three parameters increases, inference becomes slower. The dimensionality of the parameter space has the strongest impact on overall inference speed, since the autoregressive use of TabPFN necessitates the reprocessing of the adapted context for each additional parameter dimension. For the largest parameter and observation dimensions and a context of $10^{4}$ simulations, inference can take several minutes. However, this can be alleviated by filtering; for example, by selecting the $10^{3}$ most relevant simulations (see Sec. D.6 for details).
Table D-2: Inference speed of NPE-PFN. Inference speed to sample a batch of $10^{4}$ samples from the posterior measured in seconds (on an Nvidia A100 GPU). Rows show increasing dimensionality of the parameter space, columns show increasing dimensionality of the observation space, stratified by different simulation budgets.
$10^{2}$ Sim. $10^{3}$ Sim. $10^{4}$ Sim.
$\bm{\theta}$$\bm{x}$} 2 4 8 16 32 2 4 8 16 32 2 4 8 16 32
2 6 4 5 6 7 7 6 5 9 8 7 8 10 14 23
4 11 10 10 11 13 10 11 11 13 17 15 19 24 32 48
8 19 23 22 27 30 23 23 24 27 35 35 40 49 66 101
16 45 47 47 53 63 49 52 54 61 77 89 99 116 152 221
32 97 103 107 122 137 115 116 125 141 171 251 272 307 370 516
D.4 Embedding networks for high-dimensional data
We evaluate the performance of NPE-PFN-Infomax on high-dimensional simulation-based inference tasks, specifically focusing on a spatial SIR model [16] and an extended Lotka-Volterra task [94]. In the case of the spatial SIR task, where the true posterior is intractable, we rely on indirect metrics such as negative log-likelihood (NLL), TARP [95], and SBC for evaluation. NPE-PFN-Infomax leverages compressed embeddings learned via an infomax objective to distill key information from high-dimensional observations, which are then used as inputs to NPE-PFN.
@@ D.4 Embedding networks for high-dimensional data
Our results (Fig. D-5) indicate that such embedding networks—whether pre-trained or jointly trained—can be used to extend the applicability of NPE-PFN to high-dimensional observations. However, in most cases, this approach does not outperform an end-to-end trained NPE baseline. Interestingly, TabPFN without an embedding network (when applicable) often achieves better performance, suggesting that the compression step, independent of NPE-PFN, may be suboptimal. One possible explanation is that TabPFN, having been trained on causally structured data, can exploit underlying dependencies to improve inference accuracy. In contrast, compressing such highly structured data into low-dimensional summaries may discard critical information, limiting the effectiveness of inference.
@@
Figure D-5: High-dimensional data with embedding nets. (a) Visualization of observations in the spatial SIR task alongside posterior predictive samples obtained using NPE-PFN-Infomax. (b) Performance evaluation across high-dimensional variants of the spatial SIR task over 5 independent runs, reported in terms of the negative log-likelihood (NLL) of the true parameter and the area off the diagonal in TARP and SBC calibration analyses. (c) An instance of the Lotka-Volterra task with an extended time series (300 dimensions), evaluated using NPE-PFN and NPE-PFN-Infomax.
D.5 Comparison of autoregressive and ratio-based density evaluation
@@
In Sec. 2.4, we present a method for fast density evaluation based on density ratios, requiring only a single forward pass through the TabPFN classifier. While this approach offers a substantial computational advantage, it is inherently approximate, raising the question of how it compares to the autoregressive density evaluation. To answer this question, we compare the two methods on two tasks—the 2D Two Moons tasks and the 10D Gaussian Linear task—across varying simulation budgets. In each case, we evaluate the density of posterior samples from the trained model given an observation, using 5 random seeds. To construct the dataset for the ratio estimator, we draw $5\cdot 10^{3}$ samples each from the posterior of interest and the uniform base distribution, which is the default setting.
@@
Figure D-6: Comparison of density evaluation approaches on Two Moons and Gaussian Linear task. (a) Correlation between density estimates from the autoregressive and ratio-based approaches for two observations using $10^{4}$ simulations. (b) Pearson correlation between the two density estimates across varying simulation budgets for both tasks. (c) Evaluation time (log scale) for an increasing number of density evaluations using an Nvidia H100 GPU, showing the crossover point where the ratio-based approach becomes more efficient.
@@
We observe a strong correlation between autoregressive and ratio-based density evaluations for the Two Moons task and a somewhat weaker correlation for the higher-dimensional Gaussian Linear task (Fig. D-6a for a simulation budget of $10^{4}$). This trend persists across simulation budgets (Fig. D-6b). In terms of runtime, the ratio-based method has an initial cost due to posterior sampling, making it slower for the first evaluation. However, once samples are obtained, subsequent evaluations are significantly faster—an advantage that becomes very noticeable as the number of density evaluations increases (Fig. D-6c; note the logarithmic y-axis).
@@
Therefore, in scenarios that require repeated density evaluations for a single observation—such as rejection sampling in TSNPE-PFN—the ratio-based approach offers a highly favorable trade-off: It provides sufficient accuracy at a fraction of the computational cost.
D.6 Varying the number of filtered simulations
When filtering simulations, we always make full use of TabPFN’s recommended maximal context size of $10^{5}$ data points. However, in principle, the number of simulations $N_{\mathrm{filter}}$ selected by the filter is a hyperparameter that can be optimized. Here, we investigate the impact of this hyperparameter by evaluating the performance of NPE-PFN on the SBI benchmark tasks and a simulation budget of $10^{5}$. Specifically, we vary the number of filtered simulations $N_{\mathrm{filter}}$, setting it to 16, 64, 256, 1024, 2048, 4096, and 16384, and compare the resulting performance to our default choice of $10^{4}$.
For small values of $N_{\mathrm{filter}}$, the performance of NPE-PFN deteriorates across all tasks (Fig. D-7). In contrast, for $N_{\mathrm{filter}}=1024$ or larger, performance is very similar to that of our default choice with $10^{5}$ simulations. For some tasks, a slight decrease in performance is observed for the largest value of $N_{\mathrm{filter}}$, likely because this exceeds the recommended context size of TabPFN. Importantly, our default choice is always optimal or near-optimal. These results suggest that the recommended maximal context should be utilized fully. Nevertheless, these results also indicate that good performance can be achieved with smaller filter sizes (e.g., $N_{\mathrm{filter}}=2048$) to reduce the computational load.
Figure D-7: SBI benchmark results with varying number of filtered simulations. C2ST performance of filtered NPE-PFN as a function of the number of filtered simulations $N_{\mathrm{filter}}$. For all tasks, a simulation budget of $10^{5}$ is used. The red cross marks the default choice of $N_{\mathrm{filter}}=10^{4}$.
D.7 Varying feature and noise distributions
@@
Here, we investigate the robustness of NPE-PFN to different types of noise. To evaluate the robustness in such cases, we construct a variant of the Gaussian linear task from Lueckmann et al. [27] with non-Gaussian features and noise
@@
where $s$ is the scale parameter. We vary the feature dimension, the scale of feature and noise distributions, and the number of training simulations, and compare the performance of NPE-PFN against NPE.
We first evaluate performance via C2ST across all combinations of feature and noise distributions for $s_{\text{features}}=s_{\text{noise}}=1.0$, $\text{dim}({\bm{\theta}})=5$, and $10^{3}$ simulations. The heatmap shows that NPE-PFN outperforms NPE consistently, with while cells indicating training failure primarily for NPE (Fig. D-8a). For the example of Laplace feature and Cauchy noise distributions, we provide results as a function of simulation budget, confirming the robustness of NPE-PFN across distributional shifts (Fig. D-8b). Notably, while NPE achieves performance drops on several non-Gaussian configurations, NPE-PFN maintains stable performance.
Figure D-8: Robustness to feature and noise distributions. (a) Heatmaps showing C2ST performance for all combinations of feature and noise distributions with $\text{dim}({\bm{\theta}})=5$, $s=1.0$, and $10^{3}$ training simulations. White cells indicate failed training. (b) C2ST across training set sizes for an example combination of distributions. These panels are zoom-ins from (a), highlighting the stability of NPE-PFN under varying distributional assumptions.
These results suggest that, although TabPFN is pre-trained using priors based on structural causal models (SCMs) with uniform or Gaussian root noise, nonlinear transformations in the SCM induce rich marginal distributions and allow generalization to a wide range of distributions. NPE-PFN inherits this robustness, and its performance is largely invariant to different feature or noise distributions. Failures of NPE, especially with Cauchy features, are caused by instability in z-scoring, which NPE-PFN avoids through default preprocessing such as the Yeo–Johnson transform [96]. Thus, standard NPE performance could probably be improved by applying appropriate transformations. Because NPE-PFN takes advantage of the automatic preprocessing performed within TabPFN, it provides reliable performance without manual preprocessing, making it a robust and user-friendly solution in practical settings.
D.8 Order of autoregressive sampling
To sample from multi-dimensional (conditional) distributions with TabPFN, we use it in an autoregressive manner. An important question is whether the order in which we sample the dimensions matters for, e.g., the quality of the inferred posterior distributions. To investigate this question, we rerun NPE-PFN on the benchmark tasks from Sec. 3.1, but permute the order in which we sample the parameter dimensions. Two of the seven benchmark tasks are permutation invariant by design (Gaussian Mixture and Gaussian Linear), and we replace these with tasks where a distribution is constructed autoregressively. The first is a simple nonlinear task given by